OpenStudy (anonymous):

george and carmen went on a bicycle trip. they took a bus to their starting point and then biked the rest. they traveled 350 kilometers in total and they biked 80 kilometers more than they bussed let x= kilometer traveled by biked and y= kilometer traveled by bus Find how many kilometer the traveled by bike?

OpenStudy (anonymous):

let y= kilometers traveled by bus then x= kilometer traveled by bike=80+y x=80+y since x+y=350 solving these two equations.... x+y=80+y+y=350 80+2y=350 2y=270 y=135 so x=80+y=80+135=215

OpenStudy (anonymous):

thanks